OpenStudy (anonymous):

How do i simplify an expression when there is a fraction in it? For example: 3x - one over two (8x - 2) = 4

OpenStudy (phi):

can you use the equation editor ? is this \[ \frac{3x-1}{2(8x-2)} = 4 \]

OpenStudy (phi):

or \[ 3x - \frac{1}{2} \left(8x - 2\right) = 4 \]

OpenStudy (anonymous):

but how do i get the numbers off of x once i get 2x over 12 x

OpenStudy (phi):

First, what exactly is the equation ?

OpenStudy (phi):

There is an equation editor. Its button is on the lower left of the input area.

OpenStudy (phi):

if you got \[ \frac{2x}{12x} \] the x's would divide out (or "cancel"). it would simplify to \[ \frac{2x}{12x} = \frac{1}{6}\] which probably means you made a mistake ... normally the x does not cancel out when solving an equation

OpenStudy (anonymous):

i just posted the equatioin did it work?

OpenStudy (phi):

As you can see... no. It was lost in the ether

OpenStudy (anonymous):

\[3x - \frac{ 1 }{ 2 }\8x( - \2)=4\]

OpenStudy (phi):

do you mean \[ 3x - \frac{1}{2} \left(8x - 2\right) = 4 \]

OpenStudy (anonymous):

Yes

OpenStudy (phi):

in that case, notice you have a number ( the -1/2) multiplying stuff inside parens. that means you multiply -1/2 times each thing inside the parens. ( you "distribute" the -1/2) like this: \[ 3x+ - \frac{1}{2} \cdot 8x + - \frac{1}{2} \cdot-2 = 4 \] now simplify. \[ - \frac{1}{2} \cdot 8x = ?\] and \[ - \frac{1}{2} \cdot-2 = ? \]

OpenStudy (phi):

remember that \[ - \frac{1}{2} \cdot 8x \] means \[ - \frac{1}{2} \cdot 8 \cdot x \] you can just multiply the numbers. (think of 8 as 8/1) \[ - \frac{1}{2} \cdot \frac{8}{1} \cdot x \] now multiply top times top and bottom times bottom \[ \frac{-1\cdot 8 }{2\cdot 1} \cdot x \] I would divide the 2 into the 8 to simplify and get \[ \frac{-4}{1} x \] or just \[ -4x \]

OpenStudy (anonymous):

its so confusing

OpenStudy (phi):

can you do \[ - \frac{1}{2} \cdot-2 = ? \]

OpenStudy (anonymous):

\[\frac{ 2 }{ 2 }\]

OpenStudy (phi):

yes, the - times - makes it plus so you get 2/2 can you simplify 2/2 ? you can divide top and bottom by 2 and get 1/1 or just 1

OpenStudy (phi):

for the other problem \[ - \frac{1}{2} \cdot 8 \cdot x \] can you do just the numbers: \[ - \frac{1}{2} \cdot 8 = ? \]

OpenStudy (anonymous):

4x

OpenStudy (phi):

there is a - sign. you get -4x

OpenStudy (anonymous):

thats what i get wrong so much just the little things

OpenStudy (phi):

If you know that, then go carefully. you now have \[ 3x+ - \frac{1}{2} \cdot 8x + - \frac{1}{2} \cdot-2 = 4 \\ 3x -4x +1 = 4 \] on the left side you have 3 x's take away 4 x's. How many x's do you have ?

OpenStudy (anonymous):

-1x

OpenStudy (phi):

so now it is -x + 1 = 4 add -1 to both sides -x + 1 -1 = 4 -1 simplify -x +0 = 3 or -x = 3 can you finish?

OpenStudy (anonymous):

Then we can subtract 1 from both sides. Then the equation is -1x = 3

OpenStudy (anonymous):

now we divide -1 from both sides and our answer is x =3

OpenStudy (phi):

-x/ -1 = 3/-1 -x/-1 is x what is 3/-1 ?

OpenStudy (phi):

to do 3/-1 we see we have one minus sign, so the answer will be negative we do - (3/1) or -3

OpenStudy (anonymous):

ohhhh see thats what i always get wrong

OpenStudy (phi):

when multiplying or dividing, match up minus signs. each pair become a + so \[ \frac{ -3}{-1} \] has two minus signs. make them + and do + (3/1) or +3 or \[ \frac{ -3}{1} \] has one minus sign. so do - (3/1) or -3

OpenStudy (phi):

or if you had -1 * -2 * -3 *-4 you have two pair of - signs (4 all together ) so the answer is + + (1*2*3*4)

OpenStudy (phi):

If you have questions about something, try to find a video here http://www.khanacademy.org/math/algebra/introduction-to-algebra that talks about it.